OpenStudy (anonymous):

Show all work in simplifying 2x^2 + 8x - 42 over x^2 + 12x + 35 Be sure to list restrictions.

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Show all work in simplifying

OpenStudy (agreene):

\[\frac{2x^2 + 8x - 42}{x^2 + 12x + 35 }\] factor top and bottom to: \[\frac{2(x-3)(x+7)}{(x+5)(x+7)}\] cancel what you can: \[\frac{2(x-3)\cancel{(x+7)}}{(x+5)\cancel{(x+7)}}\] \[\frac{2(x-3)}{x+5}\] from there, you can expand if u want: \[\frac{2x-6}{x+5}\]

OpenStudy (anonymous):

Thank you for showing work